# Path-dependent Hamilton–Jacobi equations with $u$-dependence and time-measurable Hamiltonians

### v1

@@ 7. Optimal control with discount factors
Let $A$ be a topological space. Suppose that $A$ is the countable union of compact metrizable subsets of $A$ (cf. Chapter 21 of [CohenElliott]). Let $\mathcal{A}$ be the set of all Borel measurable functions from $[0,T]$ to $A$. Further data for our control problem are functions $f:[0,T]\times\Omega\times A\to{\mathbb{R}}^{d}$, $\lambda:[0,T]\times\Omega\times A\to{\mathbb{R}}_{+}$, $\ell:[0,T]\times\Omega\times A\to{\mathbb{R}}$, and $h:\Omega\to{\mathbb{R}}$.
@@
(i) For a.e. $t\in[0,T]$, the map $(x,y,z)\mapsto(f,\lambda,\ell)(x,a)$, $\Omega\times A\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}$, is continuous. Furthermore, $h$ is continuous.
@@
$\sup_{a\in A}\left(\left|{f(t,x,a)-f(t,\tilde{x},a))}\right|+\left|{\ell(t,x,a)-\ell(t,\tilde{x},a))}\right|+\left|{\lambda(t,x,a)-\lambda(t,\tilde{x},a)}\right|\right)$
$\qquad\leq L_{f}\,\sup_{s\leq t}\left|{x(s)-\tilde{x}(s)}\right|.$
@@
$\phi^{\prime}(t)$ $=f(t,\phi,\alpha(t))\text{ a.e.~{}on $(t_{0},T)$},\quad\phi(t)=x_{0}(t)\text{ on $[0,t_{0}]$}.$
@@
(7.3) $H(t,x,y,z):=\operatornamewithlimits{ess\,inf}_{a\in A}\left[\ell(t,x,a)+(f(t,x,a),z)-\lambda(t,x,a)\,y\right].$
@@
Suppose that Assumption 7.1 holds. One can show that $H$ defined by (7.3) satisfies Assumption 4.2 with $L_{H}=C_{f}$ and $M_{L}$ independent from $L$.
@@
For each $L\geq 0$ and $x_{0}\in\Omega$, there is a $C_{L,x_{0}}\geq 0$, such that $0\leq t_{0}\leq t_{1}\leq T$ and $x$, $\tilde{x}\in\mathcal{X}^{L}(t_{0},x_{0})$ imply $\left|{v(t_{0},x)-v(t_{1},x)}\right|\leq C_{L,x_{0}}(t_{1}-t_{0})$ and $\left|{v(t_{0},x)-v(t_{0},\tilde{x})}\right|\leq C_{L,x_{0}}\sup_{s\leq t_{0}}\left|{x(s)-\tilde{x}(s)}\right|$. Moreover, $v$ is continuous.
Proof.
Proceed as in the proof of Theorem 7.4 in [BK18JFA]. ∎
@@
Fix $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$ and $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Let $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
@@
$J^{a}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}\left[\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,\tilde{y}_{0}\right]\,{\mathrm{d}}s,$
@@
Let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$. Let $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Suppose that $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
@@
Next, let $\delta\in(0,T_{0}-s_{0}]$ and $a\in\mathcal{A}$. By Proposition 7.3,

### v2

@@ 7. Optimal control with discount factors
Let $A$ be a non-empty subset of a Polish space. Suppose that $A$ is the countable union of compact metrizable subsets of $A$ (cf. Chapter 21 of [CohenElliott]). Let $\mathcal{A}$ be the set of all Borel measurable functions from $[0,T]$ to $A$. Further data for our optimal control problem are functions $f:[0,T]\times\Omega\times A\to{\mathbb{R}}^{d}$, $\lambda:[0,T]\times\Omega\times A\to{\mathbb{R}}_{+}$, $\ell:[0,T]\times\Omega\times A\to{\mathbb{R}}$, and $h:\Omega\to{\mathbb{R}}$.
The following assumption is in force in this section.
@@
(i) For a.e. $t\in[0,T]$, the map $(x,a)\mapsto(f,\lambda,\ell)(t,x,a)$, $\Omega\times A\to{\mathbb{R}}^{d}\times{\mathbb{R}}_{+}\times{\mathbb{R}}$, is continuous.
@@
$\sup_{a\in A}\left(\left|{f(t,x,a)-f(t,\tilde{x},a)}\right|+\left|{\ell(t,x,a)-\ell(t,\tilde{x},a)}\right|+\left|{\lambda(t,x,a)-\lambda(t,\tilde{x},a)}\right|\right)$
$\qquad\leq L_{f}\,\sup_{s\leq t}\left|{x(s)-\tilde{x}(s)}\right|{\color[rgb]{0,0,0}\text{ and}}$
${\color[rgb]{0,0,0}\left|{h(x)-h(\tilde{x})}\right|\leq L_{f}\,\sup_{s\leq T}\left|{x(s)-\tilde{x}(s)}\right|}.$
@@
(7.1) $\phi^{\prime}(t)$ $=f(t,\phi,\alpha(t))\text{ a.e.~{}on $(t_{0},T)$},\quad\phi(t)=x_{0}(t)\text{ on $[0,t_{0}]$}.$
@@
(7.4) $H(t,x,y,z):={\color[rgb]{0,0,0}\inf_{a\in A}}\left[\ell(t,x,a)+(f(t,x,a),z)-\lambda(t,x,a)\,y\right].$
@@
Suppose that Assumption 7.1 holds. One can show that $H$ defined by (7.4) satisfies Assumption 4.2 with $L_{H}=C_{f}$ and $M_{L}$ independent from $L$. Note that time-measurability of $H$ follows from [AubinFrankowska_setValued, Theorem 8.2.11].
@@
For each $L\geq 0$ and $x_{0}\in\Omega$, there is a $C_{L,x_{0}}\geq 0$, such that $0\leq t_{0}\leq t_{1}\leq T$ and $x$, $\tilde{x}\in\mathcal{X}^{L}(t_{0},x_{0})$ imply $\left|{v(t_{0},x)-v(t_{1},x)}\right|\leq C_{L,x_{0}}(t_{1}-t_{0})$ and $\left|{v({\color[rgb]{0,0,0}t_{1}},x)-v({\color[rgb]{0,0,0}t_{1}},\tilde{x})}\right|\leq C_{L,x_{0}}\sup_{s\leq{\color[rgb]{0,0,0}t_{1}}}\left|{x(s)-\tilde{x}(s)}\right|$. Moreover, $v$ is continuous.
@@
Since $u$ is u.s.c. and $v$ is continuous (Proposition 7.4), the function $w$ is u.s.c. and the function $(s,x,\tilde{x})\mapsto\Upsilon(s,x)-w(s,x,\tilde{x})=v(s,\tilde{x})$ is l.s.c., i.e., we have the regularity required by Definition 5.1.
Now, fix $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$ and $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Let $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
@@
$J^{a}(s_{0}+\delta)$ $:=\int_{s_{0}}^{s_{0}+\delta}\left[\lambda^{a}(s)\,\chi^{a}(s)\,v(s_{0}+\delta,\phi^{a})-\lambda^{a}(s)\,\tilde{y}_{0}\right]\,{\mathrm{d}}s$
@@
First, note that $v$ is continuous (Proposition 7.4) and $u$ is l.s.c. Thus $w$ is u.s.c. and $(s,x,\tilde{x})\mapsto\Upsilon(s,x)-w(s,x,\tilde{x})=u(s,\tilde{x})$ is l.s.c., which is required by Definition 5.1 in order for $w$ to be a viscosity $C_{f}$-subsolution of (5.2).
Next, let $(s_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$. Let $\varphi\in\underline{\mathcal{A}}^{(C_{f})}w(s_{0},x_{0},\tilde{x}_{0})$ with corresponding time $T_{0}\in(s_{0},T]$. Suppose that $w(s_{0},x_{0},\tilde{x}_{0})>0$. Put
@@
Now, let $\delta\in(0,T_{0}-s_{0}]$ and $a\in\mathcal{A}$. By Proposition 7.3,